A certain freely falling object, released from rest, requires 1.65 s to travel the last 35.0 m before it hits the ground.

(a) Find the velocity of the object when it is 35.0 m above the ground. (Indicate the direction with the sign of your answer. Let the positive direction be upward.)
m/s

(b) Find the total distance the object travels during the fall.
m
Answer #1

S= ut +1/2gt^2

35 = u x 1.65 + 0.5 x 9.8 1.65x 1.65

u = - 13.127 m/s

V^2 - u^2 = 2gs

13.127 x13.127 - 0 = 2x 9.8 x s

S= 8.79m

Total s = 35 + 8.79= 43.79m